What is the velocity of the boat if the passenger shoots at an angle 60 with horizontal?The mass of the boat with the passengers is 100kg, mass of the bullet is 35g and its velocity is 400m/s



1
Expert's answer
2021-01-01T14:06:07-0500

"v=400 ~\frac ms,"

"M=100 ~kg,"

"m=0.035 ~kg."


"0=v_xm-u(M-m),"

"v_x=vcos60\u00b0,"

"u=\frac{vmcos60°}{M-m}=70~\frac{sm}{s}."